Perform the long division and check in your notebook. Select the correct answer.
(y3 - 125) ÷ (y - 5)

Answer by sudhanshu_kmr(1152)


(y3 - 125) = (y - 5)(y2+5y+5^2)= (y-5)(y2 + 5y + 25) ( by identity)


so,
(y3 - 125) ÷ (y - 5)
= (y2+5y+25)